# Shadows aspect of rotating black holes in the Einstein-AdS SU(N)-nonlinear sigma model

### v1

@@ I Introduction
Black holes (BHs), first predicted by general relativity (GR), are among the most fascinating objects in the universe and serve as vital testing grounds for various theories of gravity. They are particularly important in extreme conditions where both GR and quantum effects are relevant. A defining feature of BHs is their event horizon, which acts as a boundary from which no information, including light, can escape (Schwarzschild:1916uq, ). The formation of BHs is a natural consequence of GR (Penrose:1964wq, ), and they remain central to efforts aimed at reconciling GR with quantum mechanics (Hawking:1976ra, ). The no-hair theorem, a foundational principle in BH physics under GR, posits that any stationary, asymptotically flat, axially symmetric BH solution to the Einstein field equations (CarterPRD1968, ) can be fully described by three parameters: mass $M$, angular momentum $j$, and electric charge $Q$ (Israel:1967wq, ; Israel:1967za, ; Carter:1971zc, ; Hawking:1971vc, ; Robinson:1975bv, ). These parameters are encapsulated in the Kerr-Newman metric (Newman:1965my, ), which generalizes the Kerr solution Kerr:1963ud (a solution for a rotating, electrically neutral BH). The no-hair theorem is based on the uniqueness theorem of GR, which states that the Kerr-Newman metric is the sole solution to the Einstein-Maxwell equations that satisfy the conditions of stationarity and axial symmetry. For practical applications, BHs are often considered electrically neutral, as any residual charge is quickly neutralized through interactions with surrounding matter (Israel:1967wq, ; Israel:1967za, ; Carter:1971zc, ; Carter:1999mrq, ; Hawking:1971vc, ).
@@
There has been some debate about the mathematical foundation of the no-hair theorem, especially concerning the assumption of analyticity (Chrusciel:2012jk, ). The assumption that smooth, analytic metrics are accurate representations of BHs is crucial to the theorem, but it may not be universally applicable in all theoretical contexts. Modifications to GR, such as those found in metric affine gravity theories, allow for BHs that deviate from the Kerr solution. Testing the no-hair theorem thus involves exploring possible deviations from Kerr metrics, particularly in scenarios proposed by modified theories of gravity (MTG). Consequently, the Kerr metric is commonly used to model observed BHs. Disproving the existence of non-Kerr BHs requires both observational and theoretical efforts (Ryan:1995wh, ; Will:2005va, ).
@@
The study of BHs has seen significant advancements, especially following the groundbreaking detection of gravitational waves by LIGO in 2015. Additionally, the EHT has utilized the Very Long Baseline Interferometry (VLBI) technique to obtain the high angular resolution needed to image supermassive BHs. VLBI at a wavelength of 1.3 mm (230 GHz) with Earth-diameter-scale baselines is essential for resolving the shadows of M87* and Sgr A*, the two supermassive BHs with the largest apparent angular sizes (Johannsen:2012vz, ). The EHT’s development has led to the first images of BH shadows, from M87* (EventHorizonTelescope:2019dse, ; EventHorizonTelescope:2019pgp, ; EventHorizonTelescope:2019ggy, ; EventHorizonTelescope:2019jan, ; EventHorizonTelescope:2019ths, ; EventHorizonTelescope:2019uob, ) and Sgr A* (EventHorizonTelescope:2022exc, ; EventHorizonTelescope:2022urf, ; EventHorizonTelescope:2022apq, ; EventHorizonTelescope:2022wok, ; EventHorizonTelescope:2022wkp, ; EventHorizonTelescope:2022xqj, ), significantly enhancing our ability to test theories of gravity (Capozziello2024, ; CapozzielloJCAP2023, ; HuangJCAP2024, ) and demonstrating the alignment of the observed shadows with predictions for a Kerr BH in GR. In turn, investigating photon motions inside the Kerr BH is very pertinent to current findings. A specific category of photon orbits with fixed radii, known as spherical photon orbits, delineates the shadow silhouette and has astrophysical significance inside the photon orbits around a rotating BH spacetime (WilkinsPRD1972, ). The BH shadow refers to the dark region seen against the bright background of the accretion disk and surrounding space (PERLICK20221, ; ZareJCAP2024, ; MengPRD2023, ; FENG2025103075, ; MengPRD2022, ; UNIYAL2023101178, ; ZAREPLB2024, ). Photons that come extremely close to the BH, near the photon sphere, create this shadow, which is surrounded by a bright ring of light, the photon ring (Synge:1966okc, ; Bardeen:1973tla, ; Luminet:1979nyg, ; Cunningham:1973tf, ), as seen on the observer’s sky, with its outline marked by gravitationally lensed photons (Johannsen:2010ru, ). The study of BH shadows has become an essential tool for understanding the near-horizon geometry, and much research has focused on analyzing these shadows in the context of both GR (Falcke:1999pj, ; Shen:2005cw, ; Yumoto:2012kz, ; Atamurotov:2013sca, ; Abdujabbarov:2015xqa, ; Cunha:2018acu, ; Kumar:2018ple, ; Afrin:2021ggx, ; Hioki:2009na, ; Chen:2023wzv, ; Li:2024abk, ) and MTG (Amarilla:2010zq, ; Amarilla:2011fx, ; Amarilla:2013sj, ; Amir:2017slq, ; Singh:2017vfr, ; Mizuno:2018lxz, ; Allahyari:2019jqz, ; Papnoi:2014aaa, ; Kumar:2020hgm, ; Kumar:2020owy, ; Ghosh:2020spb, ; Guo:2020zmf, ; Afrin:2021wlj, ; Vagnozzi:2022moj, ; Vagnozzi:2019apd, ), including in quantum gravity inspired frameworks (Liu:2020ola, ; Brahma:2020eos, ; KumarWalia:2022ddq, ; Islam:2022wck, ; Afrin:2022ztr, ; Yang:2022btw, ). The study of BH shadows has also extended to higher-dimensional spacetimes within modified gravity theories, where extra dimensions are introduced (Papnoi:2014aaa, ; Singh:2017vfr, ; Amir:2017slq, ; Vagnozzi:2019apd, ). BH shadows thus provide a powerful tool for strong-field gravitational tests and, ultimately, for probing the validity of the no-hair theorem (Johannsen:2010ru, ; Baker:2014zba, ; Glampedakis:2023eek, ; Afrin:2021imp, ; Afrin:2021wlj, ; Afrin:2022ztr, ; Islam:2022wck, ).
@@
At the same time, the nonlinear sigma model (NLSM) has become a crucial effective field theory, both in theoretical and phenomenological terms. It has widespread applications in quantum field theory, statistical mechanics, and string theory (manton2004topological, ). One of the key uses of the NLSM is describing the low-energy dynamics of pions (nair2006quantum, ), typically in the context of the internal symmetry group $\mathrm{SU(2)}$, which represents the two-flavor case. While the predictions made by the NLSM align well with experimental results, solving the field equations derived from the model can be complex, as these equations are generally nonlinear and coupled, involving $(N^{2}-1)$ equations where $N$ is the flavor number encoded in the $\mathrm{\mathrm{SU(N)}}$ group. Consequently, many solutions have been constructed numerically. Some important results in this field are presented in Refs. (Heusler:1991xx, ; Luckock:1986tr, ; Glendenning:1988qy, ; Nelmes:2011zz, ). When the NLSM is coupled with GR or Maxwell theory to describe more intricate physical systems, the resulting field equations become even more complex. Nonetheless, in Refs. (8, ; 9, ), ansatzs have been proposed that go beyond spherical symmetry (the generalized hedgehog ansatz), which have been used to find exact solutions not only in the NLSM, but also in the Skyrme model, the generalized Skyrme model, and the Yang-Mills-Higgs theory. These solutions describe a variety of objects, including boson stars (10, ), BHs (9, ; 11, ; 12, ), black strings (14, ), gravitating solitons (15, ; 16, ), topological solitons at finite volume (17, ), and crystalline structures of topological solitons (18, ; 19, ).
@@
In this paper, motivated by the study of BHs in MTG, we consider a static spherically symmetric BH solution in an asymptotically Anti-de Sitter (AdS) spacetime, derived from the Einstein-$\mathrm{SU(N)}$-NLSM, as proposed in Ref. (Henriquez-Baez:2022ubu, ). We then apply a modified version of the Newman-Janis algorithm (NJA) NewmanJanis:1965 to obtain the rotating counterpart. The modified NJA (MNJA), as discussed in Refs. Azreg:2014 ; Azreg:2014_1 which proves particularly effective when applied to theories extending beyond GR, offering enhanced accuracy and stability in these frameworks. The primary objective of this study is to perform both analytical and numerical investigations of photon orbits and the shadow of the rotating BH spacetime, with the goal of constraining its parameters based on the observational data from the EHT.
@@
Throughout the paper, we adopt natural units with $G=c=1$, and use the (,+,+,+)(-,+,+,+)( - , + , + , + ) sign convention. Primed symbols denote derivatives with respect to the radial coordinate.
@@ II The Einstein-$\mathrm{SU(N)}$-NLSM and its AdS BH solution
An analysis of BH solutions within the framework of the Einstein-$\mathrm{SU(N)}$-NLSM is crucial for evaluating the physical viability of the model. The Einstein-$\mathrm{SU(N)}$-NLSM is characterized by the action Henriquez-Baez:2022ubu
@@
Here, $U(x)\in\mathrm{SU(N)}$, with $N$ denoting the number of flavors embedded within the $\mathrm{SU(N)}$ Lie group Bertini:2005rc ; Cacciatori:2012qi ; Tilma:2004kp , and $t_{i}$ being the generators of the $\mathrm{SU(N)}$ Lie algebra, where $i=1,\dots,(N^{2}-1)$. In this context, $\kappa$ is the gravitational constant, while $K$ is a positive coupling constant determined through experimental observations.
The Einstein-$\mathrm{SU(N)}$-NLSM framework gives rise to deformed Schwarzschild-AdS BHs Henriquez-Baez:2022ubu , describing an asymptotically AdS, static, and spherically symmetric BH spacetime. The corresponding line element is given by
@@
The notable characteristics of this class of BHs are twofold: first, they are supported by pionic matter, and second, they generalize the $\mathrm{SU(N)}$ BH analyzed in Canfora:2013osa (excluding the Skyrme term) and Gibbons:1990um . Notably, this BH solution exhibits an asymptotic structure corresponding to the AdS version of the Barriola-Vilenkin metric Rhie:1990kc and reduces to the Schwarzschild-AdS spacetime in the limit $K=0$.
@@ II.1 The RASN-BH solution
Here, $a$ denotes the BH spin parameter, while the parameters $K$, $N$, and $\kappa$, characterize the deviations of the RASN-BH from the standard Kerr-AdS BH spacetime. The coefficients of the metric (3) are independent of the time coordinate $t$ and the azimuthal angle $\phi$, implying that $\partial_{t}$ and $\partial_{\phi}$ are Killing vector fields. Consequently, any linear combination of these two Killing vector fields is also a Killing vector Visser:2007fj . Specifically, the Kerr BH is recovered in the special case where $K=\Lambda=0$, and the static metric (3) with the lapse function (4) is obtained when $a=0$. In Fig. 3, the radial profile of the $\Delta_{r}$ function has been plotted, showing its sensitivity to changes in the BH parameters.
@@ III Photon orbits and the BH shadow
Based on the four-momentum $p_{\alpha}=\partial S/\partial x_{\alpha}$, and using Carter’s separability prescription Chandrasekhar:2002 , the action can be separated as
$S=\frac{1}{2}\mu^{2}\lambda-Et+L\phi+S_{r}(r)+S_{\theta}(\theta).$ (14)
@@
Upon implementing into account the RASN-BH’s geometry, the equations of motion according to the four differential equations in the $r$-$\theta$ plane can be stated as Slany:2020jhs ; PhysRevD.81.044020 ; StuchlikEPJC2018
@@
In the above expressions, $\mathcal{O}=\mathscr{Q}-(L-aE)^{2}$ refers to the generalized Carter’s constant Carter:1968rr , where $\mathscr{Q}$ is the Carter’s separation constant. The photon trajectories around the RASN-BH are governed by the four equations (17)-(20). These trajectories depend fundamentally on the impact parameters $\xi=L/E$ and $\eta=\mathscr{Q}/E^{2}$ (Chandrasekhar:2002, ). Broadly speaking, photons exhibit three different types of trajectories: scattering, spherical, and plunging orbits. Additionally, with respect to $\xi$ and $\eta$, the radial equations (21a) and (21b) take the following forms:
@@ III.1 Orbits of constant radius and the photon regions
In fact, the study of BH shadow relies strictly on photon orbits corresponding to paths where the radial coordinate remains constant, closely skimming the BH’s horizon. These orbits, known as spherical photon orbits, play a critical role in determining whether a photon escapes to infinity or plunges into the BH. Due to their inherent instability, spherical photon orbits form an infinite sequence of photon rings that delineate the BH’s shadow. In static spacetimes, such as Schwarzschild, these orbits are confined to a single plane. However, in rotating BHs, the frame-dragging effect induces a photon region where spherical orbits become non-planar. This region, bounded by the innermost and outermost circular orbits, defines the domain of spherical photon orbits, as studied extensively in Kerr BH solutions (see Refs. Chandrasekhar:2002 ; Bardeen:1972a ; Bardeen:1973b ). Investigations into the properties and observational implications of these orbits have expanded significantly in Kerr and Kerr-like spacetimes (see Refs. stoghianidis_polar_1987 ; cramer_using_1997 ; Teo:2003 ; Johannsen:2013 ; Grenzebach:2014 ; Perlick:2017 ; charbulak_spherical_2018 ; Johnson_universal_2020 ; Himwich:2020 ; Gelles:2021 ; Ayzenberg:2022 ; Das:2022 ; fathi_spherical_2023 ; ANJUM2023101195 ; Chen:2023 ; andaru_spherical_2023 ).
Based on the formalism presented above, the spherical photon orbits at a given radius $r_{p}$ are determined by the criteria outlined in Teo:2003 :
@@
It is important to note that for a distant observer, the photons residing within the photon regions (i.e., photons on spherical orbits) form the innermost photon rings, as they complete numerous half-orbits around the BH before reaching the observer (for further discussion, see Refs. Gralla:2019 , bisnovatyi-kogan_analytical_2022 , tsupko_shape_2022 , as well as the foundational works in Refs. claudel_geometry_2001 , virbhadra_relativistic_2009 , virbhadra_compactness_2024 ). These photons thus play a key role in defining the true boundary of the shadow (or the critical curve), which will be explored in the next subsection.

### v2

@@ I Introduction
Black holes (BHs), first predicted by general relativity (GR), are among the most fascinating objects in the universe and serve as vital testing grounds for various theories of gravity. They are particularly important in extreme conditions where both GR and quantum effects are relevant. A defining feature of BHs is their event horizon, which acts as a boundary from which no information, including light, can escape [1]. The formation of BHs is a natural consequence of GR [2], and they remain central to efforts aimed at reconciling GR with quantum mechanics [3]. The no-hair theorem, a foundational principle in BH physics under GR, posits that any stationary, asymptotically flat, axially symmetric BH solution to the Einstein field equations [4] can be fully described by three parameters: mass $M$, angular momentum $j$, and electric charge $Q$ [5, 6, 7, 8, 9]. These parameters are encapsulated in the Kerr-Newman metric [10], which generalizes the Kerr solution [11] (a solution for a rotating, electrically neutral BH). The no-hair theorem is based on the uniqueness theorem of GR, which states that the Kerr-Newman metric is the sole solution to the Einstein-Maxwell equations that satisfy the conditions of stationarity and axial symmetry. For practical applications, BHs are often considered electrically neutral, as any residual charge is quickly neutralized through interactions with surrounding matter [5, 6, 7, 12, 8].
@@
There has been some debate about the mathematical foundation of the no-hair theorem, especially concerning the assumption of analyticity [13]. The assumption that smooth, analytic metrics are accurate representations of BHs is crucial to the theorem, but it may not be universally applicable in all theoretical contexts. Modifications to GR, such as those found in metric affine gravity theories, allow for BHs that deviate from the Kerr solution. Testing the no-hair theorem thus involves exploring possible deviations from Kerr metrics, particularly in scenarios proposed by modified theories of gravity (MTG). Consequently, the Kerr metric is commonly used to model observed BHs. Disproving the existence of non-Kerr BHs requires both observational and theoretical efforts [14, 15].
@@
The study of BHs has seen significant advancements, especially following the groundbreaking detection of gravitational waves by LIGO in 2015. Additionally, the EHT has utilized the Very Long Baseline Interferometry (VLBI) technique to obtain the high angular resolution needed to image supermassive BHs. VLBI at a wavelength of 1.3 mm (230 GHz) with Earth-diameter-scale baselines is essential for resolving the shadows of M87* and Sgr A*, the two supermassive BHs with the largest apparent angular sizes [16]. The EHT’s development has led to the first images of BH shadows, from M87* [17, 18, 19, 20, 21, 22] and Sgr A* [23, 24, 25, 26, 27, 28], significantly enhancing our ability to test theories of gravity [29, 30, 31, 32] and demonstrating the alignment of the observed shadows with predictions for a Kerr BH in GR. In turn, investigating photon motions inside the Kerr BH is very pertinent to current findings. A specific category of photon orbits with fixed radii, known as spherical photon orbits, delineates the shadow silhouette and has astrophysical significance inside the photon orbits around a rotating BH spacetime [33]. The BH shadow refers to the dark region seen against the bright background of the accretion disk and surrounding space [34, 35, 36, 37, 38, 39, 40]. Photons that come extremely close to the BH, near the photon sphere, create this shadow, which is surrounded by a bright ring of light, the photon ring [41, 42, 43, 44], as seen on the observer’s sky, with its outline marked by gravitationally lensed photons [45]. The study of BH shadows has become an essential tool for understanding the near-horizon geometry, and much research has focused on analyzing these shadows in the context of both GR [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59] and MTG [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79], including in quantum gravity inspired frameworks [80, 81, 82, 83, 84, 85, 86, 87]. The study of BH shadows has also extended to higher-dimensional spacetimes within modified gravity theories, where extra dimensions are introduced [63, 65, 64, 68, 88, 89]. BH shadows thus provide a powerful tool for strong-field gravitational tests and, ultimately, for probing the validity of the no-hair theorem [45, 90, 91, 92, 74, 84, 83].
@@
At the same time, the nonlinear sigma model (NLSM) has become a crucial effective field theory, both in theoretical and phenomenological terms. It has widespread applications in quantum field theory, statistical mechanics, and string theory [93]. One of the key uses of the NLSM is describing the low-energy dynamics of pions [94], typically in the context of the internal symmetry group $\mathrm{SU(2)}$, which represents the two-flavor case. While the predictions made by the NLSM align well with experimental results, solving the field equations derived from the model can be complex, as these equations are generally nonlinear and coupled, involving $(N^{2}-1)$ equations where $N$ is the flavor number encoded in the $\mathrm{\mathrm{SU}(N)}$ group. Consequently, many solutions have been constructed numerically. Some important results in this field are presented in Refs. [95, 96, 97, 98]. When the NLSM is coupled with GR or Maxwell theory to describe more intricate physical systems, the resulting field equations become even more complex. Nonetheless, in Refs. [99, 100], ansatzs have been proposed that go beyond spherical symmetry (the generalized hedgehog ansatz), which have been used to find exact solutions not only in the NLSM, but also in the Skyrme model, the generalized Skyrme model, and the Yang-Mills-Higgs theory. These solutions describe a variety of objects, including boson stars [101], BHs [100, 102, 103], black strings [104], gravitating solitons [105, 106], topological solitons at finite volume [107], and crystalline structures of topological solitons [108, 109].
@@
In this paper, motivated by the study of BHs in MTG, we consider a static spherically symmetric BH solution in an asymptotically Anti-de Sitter (AdS) spacetime, derived from the Einstein-$\mathrm{SU}(N)$-NLSM, as proposed in Ref. [110]. We then apply a modified version of the Newman-Janis algorithm (NJA) [111] to obtain the rotating counterpart. The modified NJA (MNJA), as discussed in Refs. [112, 113] which proves particularly effective when applied to theories extending beyond GR, offering enhanced accuracy and stability in these frameworks. The primary objective of this study is to perform both analytical and numerical investigations of photon orbits and the shadow of the rotating BH spacetime, with the goal of constraining its parameters based on the observational data from the EHT.
@@
Throughout the paper, we adopt geometric units with $G=c=1$, and use the (,+,+,+)(-,+,+,+)( - , + , + , + ) sign convention. Primed symbols denote derivatives with respect to the radial coordinate.
@@ II The Einstein-$\mathrm{SU}(N)$-NLSM and its AdS BH solution
An analysis of BH solutions within the framework of the Einstein-$\mathrm{SU}(N)$-NLSM is crucial for evaluating the physical viability of the model. The Einstein-$\mathrm{SU}(N)$-NLSM is characterized by the action [110]
@@
Here, $U(x)\in\mathrm{SU}(N)$, with $N$ denoting the number of flavors embedded within the $\mathrm{SU}(N)$ Lie group [114, 115, 116], and $t_{i}$ being the generators of the $\mathrm{SU}(N)$ Lie algebra, where $i=1,\dots,(N^{2}-1)$. In this context, $\kappa$ is the gravitational constant, while $K$ is a positive coupling constant determined through experimental observations.
The Einstein-$\mathrm{SU}(N)$-NLSM framework gives rise to deformed Schwarzschild-AdS BHs [110], describing an asymptotically AdS, static, and spherically symmetric BH spacetime. The corresponding line element is given by
@@
One important property of the lapse function (4) is the presence of a deficit angle, which is explicitly manifested in the term $K\kappa a_{N}$. This characteristic is commonly observed in black hole spacetimes associated with a cloud of strings (see, for instance, Refs. [117, 118, 119]). It has been demonstrated that the deficit angle can influence the motion of both massive and massless test particles by appearing in the equations of motion (see, e.g., Refs. [120, 121, 122]).
The notable characteristics of this class of BHs are twofold: first, they are supported by pionic matter, and second, they generalize the $\mathrm{SU}(N)$ BH analyzed in [123] (excluding the Skyrme term) and [124]. Notably, this BH solution exhibits an asymptotic structure corresponding to the AdS version of the Barriola-Vilenkin metric [125] and reduces to the Schwarzschild-AdS spacetime in the limit $K=0$.
@@ II.1 The RASN-BH solution
Here, $a$ denotes the BH spin parameter, while The parameters $K$, $N$, and $\kappa$ characterize the deviations of the RASN-BH from the standard Kerr-AdS BH spacetime and are directly linked to the deficit angle present in the seed metric function in Eq. (4). As analyzed in this paper, each of these parameters has distinct effects on the spacetime structure, leading to modifications in the shape of the black hole shadow. These effects will be discussed in detail in the following sections.
The coefficients of the metric (3) are independent of the time coordinate $t$ and the azimuthal angle $\phi$, implying that $\partial_{t}$ and $\partial_{\phi}$ are Killing vector fields. Consequently, any linear combination of these two Killing vector fields is also a Killing vector [126]. Specifically, the Kerr BH is recovered in the special case where $K=\Lambda=0$, and the static metric (3) with the lapse function (4) is obtained when $a=0$. In Fig. 3, the radial profile of the $\Delta_{r}$ function has been plotted, showing its sensitivity to changes in the BH parameters.
@@ III Photon orbits and the BH shadow
Based on the four-momentum $p_{\alpha}=\partial S/\partial x_{\alpha}$, and using Carter’s separability prescription [127], the action can be separated as
$S=\frac{1}{2}\mu^{2}\lambda-Et+L\phi+S_{r}(r)+S_{\theta}(\theta),$ (14)
@@
Upon implementing into account the RASN-BH’s geometry, the equations of motion according to the four differential equations in the $r$-$\theta$ plane can be stated as [128, 129, 130]
@@
In the above expressions, $\mathcal{O}=\mathscr{Q}-(L-aE)^{2}$ refers to the generalized Carter’s constant [131], where $\mathscr{Q}$ is the Carter’s separation constant. The photon trajectories around the RASN-BH are governed by the four equations (17)-(20). These trajectories depend fundamentally on the impact parameters $\xi=L/E$ and $\eta=\mathscr{Q}/E^{2}$ [127]. Broadly speaking, photons exhibit three different types of trajectories: scattering, spherical, and plunging orbits. Additionally, with respect to $\xi$ and $\eta$, the radial equations (21a) and (21b) take the following forms:
@@ III.1 Orbits of constant radius and the photon regions
In fact, the study of BH shadow relies strictly on photon orbits corresponding to paths where the radial coordinate remains constant, closely skimming the BH’s horizon. These orbits, known as spherical photon orbits, play a critical role in determining whether a photon escapes to infinity or plunges into the BH. Due to their inherent instability, spherical photon orbits form an infinite sequence of photon rings that delineate the BH’s shadow. In static spacetimes, such as Schwarzschild, these orbits are confined to a single plane. However, in rotating BHs, the frame-dragging effect induces a photon region where spherical orbits become non-planar. This region, bounded by the innermost and outermost circular orbits, defines the domain of spherical photon orbits, as studied extensively in Kerr BH solutions (see Refs. [127, 132, 133]). Investigations into the properties and observational implications of these orbits have expanded significantly in Kerr and Kerr-like spacetimes (see Refs. [134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149]).
Based on the formalism presented above, the spherical photon orbits at a given radius $r_{p}$ are determined by the criteria outlined in [136]:
@@
It is important to note that for a distant observer, the photons residing within the photon regions (i.e., photons on spherical orbits) form the innermost photon rings, as they complete numerous half-orbits around the BH before reaching the observer (for further discussion, see Refs. [150], [151], [152], as well as the foundational works in Refs. [153], [154], [155]). These photons thus play a key role in defining the true boundary of the shadow (or the critical curve), which will be explored in the next subsection.